OpenStudy (tiffany_rhodes):

Let X be a set of real numbers with least upper bound a. Prove that if epsilon > 0, then there exists an x in X, such that (a - epsilon) < x <= a.

OpenStudy (anonymous):

work directly from the definition of least upper bound that is the point of the question if what you had was not true (so i guess you are working by contradiction) then \(a\) would not be a least upper bound

OpenStudy (anonymous):

in other words, if there was no such \(x\) with \(a-\epsilon<x\leq a\) then \(a\) would not be a LEAST upper bound

OpenStudy (anonymous):

if that is not clear, let me know

OpenStudy (tiffany_rhodes):

Yeah, I guess I'm having trouble where to get started. For proof by contradiction would I assume that epsilon is less than 0? Or would I assume x >= a. I'm still kind of processing the definition of LUB.

OpenStudy (tiffany_rhodes):

*trouble understanding where

OpenStudy (tiffany_rhodes):

I don't quite understand where epsilon comes in.

OpenStudy (anonymous):

ok i can see why this would be confusing so lets go slow

OpenStudy (anonymous):

the definition of least upper bound, at least i am assuming this is the definition you are using is this: \(a\) is a least upper bound of a set of real numbers \(X\) if i) for every \(x\in X, x\le

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

i) for every \(x\in X, x\leq a\) and ii) if \(k\) is another upper bound for \(X\) then \(k\geq a\) is that what you are using?

OpenStudy (tiffany_rhodes):

yes, those definitions are in my textbook

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

so lets work by contradiction and suppose that the statement above is not true, i.e it is not true that given any \(\epsilon>0\) there exists an \(x\in X\) such that \(a-\epsilon <x\leq a\)

OpenStudy (anonymous):

do you know how to negate that statement? what it means for it not to be true? "no" is a fine answer, just asking

OpenStudy (tiffany_rhodes):

yes, wouldn't it be for all x in X, ( a- epsilon < x <= a) is not true and then you'd just flip the inequalities?

OpenStudy (anonymous):

oh no so lets take this part and make it clear

OpenStudy (tiffany_rhodes):

oh lol, sorry. It's been awhile since my first proof class.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

to negate that statement you do not flip any inequalities suppose i said "for all houses in this neighbourhood, there is a room that is painted red" what would we do to show that it is not true? answer: find SOME house with NO red room

OpenStudy (anonymous):

the negation of the statement "for every \(\epsilon>0\) there is some \(x\in X\) such that is There is some \(\epsilon >0\) such that there is NO \(x\in X\) with \(a-\epsilon <x\leq a\)

OpenStudy (anonymous):

let me know if that is clear nor not

OpenStudy (tiffany_rhodes):

okay, got it so far.

OpenStudy (anonymous):

ok so lets assume by contradtction that the statement is not true, i.e. there is some \(\epsilon>0\) with no \(x\in X\) satisfying \(a-\epsilon<x\leq a\) what does that mean about \(a-\spsilon\)?

OpenStudy (anonymous):

oops about \(a-\epsilon\)?

OpenStudy (anonymous):

it means i) \(a-\epsilon \) is an upper bound for \(X\) since there are no \(x\)'s larger than it

OpenStudy (anonymous):

and also since \(a-\epsilon <a\) that means \(a\) is NOT the least upper bound

OpenStudy (anonymous):

i.e. it contradicts the fact that \(a\) was a least upper bound since we have another upper bound that is less than it

OpenStudy (tiffany_rhodes):

which would be a contradiction because the problem states that the LUB for X is a?

OpenStudy (anonymous):

yes, exactly

OpenStudy (anonymous):

we have found another upper bound, namely \(a-\epsilon\) that is clearly less than \(a\)

OpenStudy (tiffany_rhodes):

oh okay. and that's where that 2 condition of LUB comes in?

OpenStudy (tiffany_rhodes):

oh never mind. I'm referring to my textbook

OpenStudy (anonymous):

yes, \(a\) was not only supposed to be the upper bound but also if \(k\) is some other upper bound then \(k\geq a\)

OpenStudy (anonymous):

i hope the logic of that is more or less clear

OpenStudy (tiffany_rhodes):

Yeah, the logic is completely clear. The most difficult part is really understanding the definition of LUB and the role epsilon plays even though I know what a LUB is intuitively

OpenStudy (tiffany_rhodes):

thanks for your help :)

OpenStudy (anonymous):

yeah don't get messed up by that greek letter, it just means some positive number

OpenStudy (anonymous):

you're welcome